OpenStudy (anonymous):

What algebraic property does this statement show? 3 + (7 + 2) = (3 + 7) + 2 associative property closure property commutative property symmetric property

OpenStudy (anonymous):

The basic algebraic properties of real numbers a,b and c are: 1. Closure: a + b and ab are real numbers 2. Commutative: a + b = b + a, ab = ba 3. Associative: (a+b) + c = a + (b+c), (ab)c = a(bc) 4. Distributive: (a+b)c = ac+bc 5. Identity: a+0 = 0+a = a 6. Inverse: a + (-a) = 0, a(1/a) = 1 7. Cancelation: If a+x=a+y, then x=y 8. Zero-factor: a0 = 0a = 0 9. Negation: -(-a) = a, (-a)b= a(-b) = -(ab), (-a)(-b) = ab

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

Cant u read?

OpenStudy (anonymous):

. Associative: (a+b) + c = a + (b+c), (ab)c = a(bc)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yw